ABC has completed the following inventory purchases budget. January Inventory 12,000 purchases budget in units Each unit costJanuary February Inventory 12,000 15,000 purchases budget in units Each unit costs $5. ABC pays 20% in the month of purchaseJanuary February Inventory 12,000 15,000 purchases budget in units Each unit costs $5. ABC pays 20% in the month of purchaseJanuary February Inventory 12,000 15,000 purchases budget in units Each unit costs $5. ABC pays 20% in the month of purchase
Answer #1

Answers

  • Answer #1
    >Cash disbursed in January = 12000 units x $ 5 = $ 60,000
    Correct Answer = Option #3: $ 60,000
  • Answer #2
    >Cash disbursed in January = ($25000) + (12000units x $5 x 20%)
    = 25000 + 12000
    = $ 37,000
    Correct Answer = Option #4: $ 37,000
  • Answer #3
    >Cash disbursed in January = ($70000 x 80%) + (12000units x $5 x 20%)
    = 56000 + 12000
    = $ 68000
    Correct Answer = Option #2: $ 68000
  • Answer #4
    >Cash disbursed in Feb = (12000units x $5 x 80%) + (15000 units x $ 5 x 20%)
    = 48000 + 15000
    = $ 63000
    Correct Answer = Option #2: $ 63000